OpenStudy (vera_ewing):

A gas has a volume of 4.25 m3 at a temperature of 95.0°C and a pressure of 1.05 atm. What temperature will the gas have at a pressure of 1.58 atm and a volume of 2.46 m3? A. 356°C B. 321°C C. 47.5°C D. 82.7°C

OpenStudy (vera_ewing):

@sweetburger

OpenStudy (vera_ewing):

@Abmon98

OpenStudy (vera_ewing):

@aaronq

OpenStudy (sweetburger):

so the equation to be used is PV/T = PV/T

OpenStudy (vera_ewing):

Ok the answer is D right?

OpenStudy (sweetburger):

i got an answer of 47.7 closer to C.

OpenStudy (vera_ewing):

Ok thanks. :)

OpenStudy (sweetburger):

(4.25 x 1.05) / (368) = (1.58 x 2.46)/X X= 320.5K then i subtracted 273 and got 47.5